3. In each of the following cases explain whether R2"\to" R is a linear transformation, if it is, supply a proof, if not, supply a counter ample

a) T(a, b) =a + b

b) T(a, b) =ab

c) T(a, b) =|a|2

d) T(a, b) =a - b



1
Expert's answer
2021-11-09T16:52:07-0500

"\begin{aligned}T: \mathbb{R^2} \rightarrow \mathbb{R}\\ (a)\ T(a, b)=a+b & \\ T\left(\left(a_{1}, b_{1}\right)+\left(a_{2}, b_{2}\right)\right) &=T\left(a_{1}+a_{2}, b_{1}+b_{2}\right) \\ &=a_{1}+a_{2}+b_{1}+b_{2} \\ &=a_{1}+b_{1}+a_{2}+b_{2} \\ &=T\left(a_{1}, b_{1}\right)+T\left(a_{2}, b_{2}\right) \end{aligned}"

"\begin{aligned} T(\alpha(a, b))=& T(\alpha a, \alpha b) \\ &=\alpha a+\alpha b=\alpha(a+b)=\alpha T(a, b) \end{aligned}"

Thus "T" is linear transformation.


"(b) \quad T(a, b)=a b\\
\text{ This is not linear transformation}"

"\begin{aligned}
&\text { as } T(1,1)=1, T(1,0)+T(0,1)=0+0=0 \\
&\Rightarrow T(1,1) \neq T(1,0)+T(0,1)
\end{aligned}"


"(c)\ T(a, b)=|a|^{b}\\
\text{This is not linear transformation.}"

"\begin{aligned}
&T(2,1)=2^{1}=2 \\
&T(2,0)=2^{0}=1 \quad \& \quad T(0,1)=0^{1}=0 \\
&T(2,0)+T(0,1)=1 \neq T(2,1)
\end{aligned}"


(d) "T(a, b)=a-b"

"\begin{aligned} & \\ T\left(\alpha\left(a_{1}, b_{1}\right)+\left(a_{2}, b_{2}\right)\right) &=T\left(\alpha a_{1}+a_{2}, \alpha b_{1}+b_{2}\right) \\ &=\alpha a_{1}+a_{2}-\alpha b_{1}-b_{2} \\ &=\alpha\left(a_{1}-b_{1}\right)+a_{2}-b_{2} \\ &=\alpha T\left(a_{1}, b_{1}\right)+T\left(a_{2}, b_{2}\right) \end{aligned}"

Thus, T is linear transformation.